OpenStudy (ammarah):

Geometry Questions!! Whose smarter than a 5th grader??? Help!! Ill give u a medal!!!!! =)

OpenStudy (ammarah):

Find the scale factor of ABCD to KLMN

OpenStudy (ammarah):



OpenStudy (raden):

the second figure is cut, i didnt see it

OpenStudy (anonymous):

Fifth graders don't learn this >.>

OpenStudy (ammarah):

i know i just wanted ppl to come help me

OpenStudy (mathstudent55):

Since you know the lengths of two corresponding sides in the two quadrilaterals, you can calculate the scale. Side AD corresponds to side KN. AD = 60 and KN = 36, so AD/KN = 60/36 =5/3

OpenStudy (ammarah):

thanks

OpenStudy (anonymous):

How do you know that the quadrilaterals are similar?

OpenStudy (ammarah):

Someone can read it

OpenStudy (ammarah):

cuz there similar...

OpenStudy (ammarah):

how old r u grazes?

OpenStudy (ammarah):

??? cat got ur tongue?

OpenStudy (mathstudent55):

Grazes' question is valid. The scale factor only applies if the two quadrilaterals are similar.

OpenStudy (ammarah):

i know they are similar i already solved the problem..someone else helped me on it....... thanks u guys anyways....

OpenStudy (anonymous):

Even if a cat had caught my tongue, I would still be able to type.

OpenStudy (ammarah):

lol i know thats why i said it.........i was going to say cat gor ur hand!!lol

OpenStudy (anonymous):

Which is why I have two hands. To prevent the occurrence of hand-grabbing cats.

OpenStudy (ammarah):

cat got both of ur hands at once!! or you have two cats in the house!! =)

OpenStudy (ammarah):

how old r u? btw

OpenStudy (ammarah):

and r u f or m?

OpenStudy (anonymous):

15m.... O_o